OpenStudy (helpmeee123):

Can someone help me with this plz? Plz show ur work .... A local high school collected $1590 from 321 people who attended a football game. The price of each adult admission is $6. People between the ages of 4-17 paid a children admission rate of $4. How many adult tickets and child tickets were sold that day? Please help me with the elimination thing. Thankyou

OpenStudy (michele_laino):

Please note that your problem, can be modeled by the subsequent algebraic system: \[\left\{ \begin{gathered} 6A + 4C = 1590 \hfill \\ A + C = 321 \hfill \\ \end{gathered} \right.\] where A is the cost of each adult ticket, and C is the cost of each children ticket